A brass rod is .500 m long at 20.0°C. What is the length of the rod if it is heated
to 50.0°C?
1
Expert's answer
2020-04-20T10:43:14-0400

The linear thermal expansion is given by

"l=l_0(1+\alpha \Delta T)"

"l=0.500(1+19\times 10^{-6}\times 30)=0.5003 \:\rm m"